sqrt7a+6sqrt6^3a^3=
Answer by dabanfield(803)

sqrt7a+6sqrt6^3a^3=
sqrt(7a) + 6*(sqrt(6*6^2*a*a^2) =
sqrt(7a) + 6*sqrt(6)*sqrt(6^2)*sqrt(a)*sqrt(a^2) =
sqrt(7a) + 6*sqrt(6)*6*sqrt(a)*a =
sqrt(7a) + (36a)*sqrt(6)*sqrt(a) =
sqrt(7a) + (36a)sqrt(6)sqrt(a)